OpenStudy (anonymous):

How can -4x^2-6x+4 be factored?

OpenStudy (anonymous):

what method u need to use?

OpenStudy (anonymous):

I'm supposed to solve it. \[\sqrt{4-6x}=2x\] is the original

OpenStudy (anonymous):

I know you square both sides like so \[(\sqrt{4-6x})^2=(2x)^2\] Which get you \[4-6x=4x^2\] I think.

OpenStudy (anonymous):

divide by 2 \[2x^2+3x -2=0\]

OpenStudy (anonymous):

divide again by 2 \[x^2 +\frac{ 3 }{ 2 }x-1=0\]

OpenStudy (anonymous):

(x )(x ) two numbers the sum would be 3/2 the multiplication would be 1

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

\[(x+2)(x-\frac{ 1 }{ 2 })=0\]

OpenStudy (anonymous):

Thank You. That cleared it up alot.